OpenStudy (anonymous):

-8/7 divided by 64/7 =

OpenStudy (anonymous):

(-8/7)/(64/7)=(-8/7)*(7/64)=(-56/448)

OpenStudy (anonymous):

=-1/8

OpenStudy (anonymous):

I forgot my negitive when I posted my answer now I have another problem! ehh.. stupid calculators lol how aboue this one...